# The Most Common Habitable Planets III – Modeling Temperature Forcing and Surface Conditions on Rocky Exoplanets and Exomoons

### v1

@@ 7 Conclusions
In this work we improve on the model of 59, originally developed for the study of radiative-forced temperature variations on exoplanets, extending the formalism (Sec. 2) and exploring a wider range of cases for albedo, orbital eccentricity and atmospheric masses. We apply the model to estimate the brightness temperature modulation of both real and hypothetical rocky exoplanets and large exomoons, orbiting real and hypothetical host stars and host giant exoplanets inside the HZ. The results show that for specific cases an atmosphere may distort the exoplanet’s or exomoon’s profile of brightness temperature variations and serve as a measurable indication of the presence of an atmosphere.
@@
Such temperature variations do not significantly affect the star-moon contrast in these systems, which are around 10-9-10-8. In all cases studied by us, the temperature variations have too low a flux modulation effect to be detected by current instruments. Nevertheless, in the future, new instruments on ELT-class telescopes, like GMTNIRS/GMT 32, MODHIS/TMT 71 and HIRES/ELT 48, along with new space telescopes using IR nulling interferometry and upgrades to existing instruments at Keck, VLT and Subaru will provide the needed combination of high-sensitivity and high-angular resolution observations. Instruments such as ELT/METIS 5, combining coronagraphy for high contrast IR imaging and spectroscopy, will be able to perform direct characterization of exoplanets and possibly exomoons, finally bringing into reach the observations of temperature variations like those proposed by our work.
@@ Acknowledgements
The authors thank the anonymous referee for the enriching comments and suggestions. R.G.G.F. thanks CNPq/Brazil for the PIBIC undergraduate fellowship. G.F.P.M. acknowledges grant 474972/2009-7 from CNPq/Brazil. M.B.F. acknowledges financial support from CNPq/Brazil (grant number: 307711/2022-6). We thank Dr. Adrián Rodríguez for valuable discussions on tidal dissipation.

### v2

@@ 7 Conclusions
In this work we improve on the model of 57, originally developed for the study of radiative-forced temperature variations on exoplanets, extending the formalism (Sec. 2) and exploring a wider range of cases for albedo, orbital eccentricity and atmospheric masses. We apply the model to estimate the brightness temperature modulation of both real and hypothetical rocky exoplanets and large exomoons, orbiting real and hypothetical host stars and host giant exoplanets inside the HZ. The results show that for specific cases an atmosphere may distort the exoplanet’s or exomoon’s profile of brightness temperature variations and serve as a measurable indication of the presence of an atmosphere.
@@
Such temperature variations do not significantly affect the star-moon contrast in these systems, which are around 10-9-10-8. In all cases studied by us, the temperature variations have too low a flux modulation effect to be detected by current instruments. Nevertheless, in the future, new instruments on ELT-class telescopes, like GMTNIRS/GMT 31, MODHIS/TMT 69 and HIRES/ELT 46, along with new space telescopes using IR nulling interferometry and upgrades to existing instruments at Keck, VLT and Subaru will provide the needed combination of high-sensitivity and high-angular resolution observations. Instruments such as ELT/METIS 5, combining coronagraphy for high contrast IR imaging and spectroscopy, will be able to perform direct characterization of exoplanets and possibly exomoons, finally bringing into reach the observations of temperature variations like those proposed by our work.
@@ Acknowledgements
The authors thank the anonymous referee for the enriching comments and suggestions. RGGF thanks the Brazilian Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) for the PIBIC undergraduate fellowship. GFPM acknowledges grant 474972/2009-7 from the Brazilian Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq). MBF acknowledges financial support from he Brazilian Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) (grant number: 307711/2022-6). We thank Dr. Adrián Rodríguez for valuable discussions on tidal dissipation. The authors acknowledge the use of data from NASA’s planetary fact sheet webpage, produced by Dr. David R. Williams. This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program.